What is the width if the length is 95 and perimeter is 338
Answer by Theo(13342)

perimeter is equal to 2 * length + 2 * width
2 * length = 2 * 95 = 190

338 - 190 = 148

148 / 2 = 74

length is 74 and width is 95

2 * length + 2 * width = 190 + 148 = 338

this all stems from the following equation:

P = 2L + 2W

P = perimeters
L = length
W = wisth.

since you know the length and the perimeters, you want tos solve this equation for width.

you will get W = (P - 2L) / 2

since P = 338 and L = 95, this equation becomes:

W = (338 - 2 * 95) / 2

simplify to get:

W = (338 - 190) / 2

simplify further to get:

W = 148/2

solve for W to get W = 74